OpenStudy (anonymous):

In the fraction, assume that no denominator is 0. Simplify the expression. (a^2b^3)^3/a^6b^6

jimthompson5910 (jim_thompson5910):

(a^2b^3)^3 = (a^2)^3*(b^3)^3 = a^(2*3)*b^(3*3) = a^6b^9 So (a^2b^3)^3 = a^6b^9 (this is just the numerator) ----------------------------- (a^2b^3)^3/(a^6b^6) (a^6b^9)/(a^6b^6) ... simplify the numerator (see above) Now subtract corresponding exponents a^(6-6)b^(9-6) a^0b^3 b^3 So (a^2b^3)^3/(a^6b^6) = b^3

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np